# Light-induced torque in ferromagnetic metals via orbital angular momentum generated by photon-helicity

### v1

@@ Abstract
This work was partially supported by Grants-in-Aid for Scientific Research (Nos. 21H04648 and 21H05000), JST PRESTO (No. JPMJPR22B2), X-NICS of MEXT (No. JPJ011438), CSIS cooperative research project in Tohoku University, the Asahi Glass Foundation, the Murata Science Foundation, Advanced Technology Institute Research Grants, FRIS Creative Interdisciplinary Collaboration Program in Tohoku University. This work was supported by the ANR-20-CE09-0013 UFO, by “Lorraine Université d’Excellence” reference ANR-15-IDEX-04-LUE, and by the French National Research Agency through the France 2030 government grants PEPR Electronic EMCOM (ANR-22-PEEL-0009). S. I. thanks to TI-FRIS fellowship at Tohoku University. K. N. and K. I. thank to GP-Spin at Tohoku University. S. M. thanks to CSRN of CSIS at Tohoku University. The authors thank Gyung-Min Choi for valuable discussion.

### v2

@@ Abstract
This work was supported by Grant-in-Aid for Scientific Research (Nos. 21H04648 and 21H05000), Grant-in-Aid for Challenging Research (No. 24K21234), Grant-in-Aid for Transformative Research Areas (No. 24H02235), JST PRESTO (No. JPMJPR22B2), X-NICS of MEXT (No. JPJ011438), CSIS cooperative research project in Tohoku University, the Asahi Glass Foundation, the Murata Science Foundation, Advanced Technology Institute Research Grants, FRIS Creative Interdisciplinary Collaboration Program in Tohoku University. This work was also supported by the ANR-20-CE09-0013 UFO, by “Lorraine Université d’Excellence” reference ANR-15-IDEX-04-LUE, and by the French National Research Agency through the France 2030 government grants PEPR Electronic EMCOM (ANR-22-PEEL-0009). S. I. thanks to TI-FRIS fellowship at Tohoku University. K. N. and K. I. thank to GP-Spin at Tohoku University. S. M. thanks to CSRN of CSIS at Tohoku University. The authors thank Gyung-Min Choi for valuable discussion.